Grade upto college level Algebra

The sum of the first n terms of the series 12 + 2.22 +32 + 2.4 + 52 + 2.62 +is
n(n+1)2 / 2, when n is even. When n is odd, the sum is ………….

Shane Macguire
12 Years agoGrade upto college level

3 Answers

Deepak Patra
12 Years ago
Hello Student,
Please find the answer to your question
Ans. When n is odd, let n = 2m +1
∴ The req. sum
= 12 + 2.22 + 32 + 2.42 + ….. + 2(2m) 2 + (2m + 1) 2
= Σ (2m + 1) 2 + 4[12 + 22 + 32 + …. +m2 ]
=(2m + 1) (2m + 2) (4m + 2 + 1)/6 + 4m (m + 1) (2m + 1)/6
= (2m + 1) (m + 1)/6 [2 (4m + 3) + 4m ]
= (2m + 1) (2m + 2)(6m + 3)/6 = ((2m + 1)2 (2m + 2)/2
= n2(n + 1)/2 [ ∵ 2m + 1 =n]
ALTERNATE SOLUTION:
∵ n is odd, last term = n2
∴ req. sum
= [12 + 2.22 + 32 + 2.42 +….. + 2(n-1) 2 ] + n2
= (n – 1)n2 /2 +n2 [ Using sum for( n – 1 ) to be even = n2 (n + 1)/2
Thanks
Deepak Patra
askIITians Faculty
Ritik
9 Years ago
Please explain these stepes. :- 12 + 2.22 + 32 + 2.42 + ….. + 2(2m) 2 + (2m + 1) 2= Σ (2m + 1) 2 + 4[12 + 22 + 32 + …. +m2 ]
ankit singh
5 Years ago
When n is odd, it means the sum will be S(n - 1) - the last odd term ie n^2. We can get the result by substituting k = n - 1 in the S(k) expression, and subtracting n ^ 2 from the value, and simplifying it. Hope this answers your question.